OpenStudy (anonymous):

Solve the system of equations. 2x + 4y = 36 10y = 5

OpenStudy (anonymous):

solve for y , in this case y=1/2 ( second equation) , later replaces in the first equation.

OpenStudy (anonymous):

multiply 2?

OpenStudy (anonymous):

in the second equation how 10y=5 , later y=5/10 =1/2 ,you have the value of y and you can replacein the first equation and get x

OpenStudy (anonymous):

im confused..can you please show me

OpenStudy (ledah):

I believe he's suggesting the substitution method.

OpenStudy (anonymous):

2x + 4y = 36 .....A 10y = 5.....B use equation B to get the value of y y=5/10 y=1/2 place the value of y in equation A 2x+4(1/2)=36 2x+2=36 2x=36-2 2x=34 x=34/2 x=17

OpenStudy (anonymous):

so you get x=17 and y=1/2 or (17,1/2)

OpenStudy (anonymous):

@coolaidd : hope you understand it :)

OpenStudy (anonymous):

:) thanks!

OpenStudy (anonymous):

u welcum

OpenStudy (anonymous):

=)